OpenStudy (anonymous):

Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? Lance saw 2 reef sharks in 6 minutes. Jasmine saw 3 sea cucumbers in 8 minutes. Robert saw 4 jellyfish in 15 minutes. Shilpa saw 3 pencil urchins in 10 minutes.

OpenStudy (anonymous):

JASMINE

OpenStudy (anonymous):

@tommyblythe56 explain

OpenStudy (uri):

Lance is 2/6=1/3=0.333 Jasmine is 3/8=0.375 Robert is 4/15=0.266 Shipa is 3/10 =0.3 So jasmine has the fastest rate.

OpenStudy (anonymous):

@uri thanks!

OpenStudy (anonymous):

srry